# Efficient Dynamic Scene Editing via 4D Gaussian-based Static-Dynamic Separation

### v1

@@ Conclusion.
In this work, we have proposed an efficient 4D dynamic scene editing framework leveraging 4D Gaussian Splatting (4DGS) and score distillation mechanism. Focusing on the static-dynamic separability of 4DGS, we demonstrated that editing only on the static canonical component and subsequently refining any resulting motion artifacts provides a significantly faster and more efficient approach for completing dynamic scene editing. We also showed that the score distillation mechanism can effectively transfer generative priors to 4D space, providing an efficient alternative to RGB loss, which requires additional 2D image updates. Qualitative and quantitative results have demonstrated that our method outperforms the baseline in both visual metrics and editing efficiency.
@@
Our method is bounded by the capabilities of IP2P; it is not suited for motion editing and requires additional segmentation pipelines for partial editing tasks. The quality of the 4D representation also imposes limitations, as it can lead to visible motion artifacts even after temporal refinement.

### v2

@@ Conclusion.
We proposed Instruct-4DGS, an efficient 4D dynamic scene editing framework leveraging 4D Gaussian Splatting (4DGS) and a score distillation mechanism. Exploiting the static-dynamic separability of 4DGS, our approach edits only static canonical components and refines motion artifacts, significantly enhancing editing speed and efficiency. Score distillation effectively transfers generative priors into 4D space, offering an efficient alternative to the conventional RGB loss, which requires updating additional 2D images. Experimental results demonstrated superior visual quality and editing efficiency compared to the baseline.
@@
Our method relies on IP2P’s capabilities, cannot directly edit motion, requires segmentation for partial edits, and may show motion artifacts due to limitations of the 4D representation, even after temporal refinement.
Acknowledgements.
This work was supported by Institute of Information & Communications Technology Planning & Evaluation (IITP) grant funded by the Korea government (MSIT) (RS-2024-00439020, Developing Sustainable, Real-Time Generative AI for Multimodal Interaction, SW Starlab).
Supplementary Material
7 Additional Qualitative Results
7.1 Results on Monocular Datasets
While 4D dynamic scene editing typically relies on multiview video datasets to sufficiently capture spatio-temporal information, we evaluate our method on the DyCheck [16] and HyperNeRF [41] datasets to explore its potential applicability to monocular video inputs. For these monocular datasets, we cannot obtain edited multiview supervision images for editing canonical 3D Gaussians. Therefore, we skip Stage 1 (described in Sec. 4.2) and only apply Stage 2, the score-based temporal refinement (described in Sec. 4.3). Figure 9 presents a comparison between Instruct 4D-to-4D (baseline) and Instruct-4DGS (ours) on the DyCheck dataset, while Fig. 10 shows qualitative results of our method on the HyperNeRF dataset. Our Instruct-4DGS produces plausible dynamic scene editing results even on monocular datasets, and we expect the performance to further improve as techniques for reconstructing 4D Gaussians from monocular videos and editing with the SDS mechanism continue to advance.
7.2 Results with Varying Camera Poses
To further assess the spatial consistency and quality of our edited 4D Gaussian representations, we render the edited dynamic scenes from the DyNeRF [28] dataset under various camera poses. As shown in Fig. 11, the results produced by our Instruct-4DGS maintain plausible geometry and appearance across different viewpoints.
Figure 9: Qualitative comparison of visual quality on the DyCheck [16] dataset (a monocular dataset): We compare our method, Instruct-4DGS (ours), with the baseline, Instruct 4D-to-4D [38] (baseline), on the mochi-high-five scene from the DyCheck dataset.
Figure 10: Qualitative results of our Instruct-4DGS on the HyperNeRF [41] dataset (a monocular dataset): We evaluate our method on the Interp_chickchicken scene from the HyperNeRF dataset.
Figure 11: Qualitative results of our Instruct-4DGS under various camera poses on the DyNeRF [28] dataset: We render the edited dynamic scene from novel camera poses to evaluate the spatial consistency of our method. Our Instruct-4DGS produces view-consistent and geometrically plausible results.
8 Full Set of Editing Instructions
Here, we provide the full set of editing instructions used for our dynamic scene editing experiments.
We used “Make the person a statue”, “Make the person a marble Roman sculpture”, and “Make the person a wood sculpture” for Tab. 1.
We used “What if it was painted by {Makoto Shinkai, Henri Matisse, Utagawa Hiroshige, Van Gogh, Edvard Munch}?”, “Make it a Fauvism painting”, “Make the person a statue”, “Make the person a marble Roman sculpture”, and “Make the person a wood sculpture” for Fig. 5.
We used “Make the person a marble Roman sculpture”, “What if it was painted by {Van Gogh, Edvard Munch}?”, “Make it a Fauvism painting”, “Make this a cozy wooden cabin bar with soft lighting and rustic decorations”, “Turn the man into a bronze sculpture”, “Add a beautiful sunset”, “Make it underwater”, “Give him a Victorian gentleman’s attire”, “Make it a Fauvism painting”, and “What if it was painted by Van Gogh?” for Fig. 6